Can you help me step by step on how to solve: -2log base 4(x)= log base 4(3)^2
Answer by lwsshak3(11628)

Can you help me step by step on how to solve:
-2log base 4(x)= log base 4(3)^2
log base 4(3)^2+2log base 4(x)=0
log base 4(9)+log base 4(x^2)=0
place under single log
log4[9*x^2]=0
convert to exponential form: base(4) raised to log of number(0)=number(9x^2)
4^0=9x^2=1
x^2=1/9
x=-1/3(reject, x>0)
or
x=1/3